ID |
Date |
Author |
Type |
Category |
Subject |
9627
|
Wed Feb 12 14:05:16 2014 |
ericq | Update | SUS | PRM Oplev Checked Out | [ericq]
Steve fixed the PRM oplev pointing. I turned on the loops and measured the OLG, then set the pitch and yaw gains such that the upper UGF was ~8Hz (motivated by Jenne's loop design in ELOG 9401)
- Pitch gain: +7
- Yaw Gain: -5
I then measured the oplev spectra of the optics as they were aligned for PRMI. (OSEMs on, oplevs on, LSC off, and ASC off)
Next, Jenne and I need to fix the ASC loop such that it properly accounts for the oplev loop.
|
9626
|
Wed Feb 12 11:57:55 2014 |
Jenne | Update | LSC | Musings on RFPDs | [Rana, Jenne, Manasa]
We looked at the I vs. Q separation in several of the Refl PDs, while driving the PRM, while the PRMI was locked on sidebands.
For REFL 55, we adjusted the demod phase to try to minimize the peak in the Q signal, and were only able to get it to be about 1/10th the size of the I peak. This is not good, since it should be more like 1/100, at least.
For both REFL 11 and REFL 165, we were able to get the Q peaks to less than 1/100 of the I peak height.
We changed the REFL55 phase from 17 to 16, and the REFL165 phase from -160.5 to -162.5.
Since we believed that we had done a good job of setting the demod phase for REFL165, we used it to also check the balance of BS/PRM for MICH locking. I drove the BS with an arbitrary number (0.5), which creates a peak in the I phase of REFL165, and then I put in a drive on the PRM and tweaked it around until that peak was minimized. I came up with the same ratio as Koji had last Friday: BS=0.5, PRM=-0.2625. (The old ratio we were using, up until ~December when we started locking MICH with the ITMs, was BS=0.5, PRM=-0.267).
Also, while we were locked using REFL55 I&Q, we noticed that the other REFL PDs had lots of broadband noise in their I signals, as if some noise in the REFL55 diode is being injected into the PRM, that we are then seeing in the other PDs.
Some checks that we need to do:
* Inject a calibration line, set all the peak heights equal, and look at the noise floors of each PD.
* Use the calibration line to calibrate the PDs (especially REFL165) into meters, so that we know that it's noise is low enough to hold the PRC through the CARM offset reduction.
* Check out the state of the transmission QPDs - what is their noise, and is it good enough to use for holding the arms after we transition from green beatnote locking? Does the whitening switching do anything? What is the state of the whitening? |
9625
|
Tue Feb 11 22:17:06 2014 |
Koji | Update | Green Locking | ALS X and Y arm restored | Nice restoration. We eventually want to make transition of the servo part from ALS to LSC model for the further handing off to the other signals.
Please proceed to it. |
9624
|
Tue Feb 11 21:22:02 2014 |
manasa | Update | Green Locking | ALS X and Y arm restored | The X and Y arms were locked successfully using ALS and the arms could be scanned and held to support IR resonance.
The same procedure as in elog 9219 was followed. In-loop noise was measured to be between 200-300 Hz rms for the lock.
ALS settings for the lock
X arm : FM 2, 3, 5, 6, 7, 8, 10 Gain = 11.0
Y arm : FM 2, 3, 5, 6, 7, 8, 10 Gain = 10.0 |
9623
|
Tue Feb 11 16:49:36 2014 |
Steve | Update | PEM | floor cleaning | Keven, Steve
The floor was cleaned under the east arm tube with hand held wet towel. We moved staff around and mopped. I did at the bottom of rack 1Y1, 2 and 3.
Last week we did the south arm tube floor.
Next week we 'd like to clean under rack 1X1,2,3, 4, 5, 6 and 7 |
9622
|
Tue Feb 11 10:44:15 2014 |
Steve | Update | VAC | RGA scan at day 6 |
|
Attachment 1: pd77m6d.png
|
|
9621
|
Mon Feb 10 22:21:55 2014 |
manasa | Update | Green Locking | X and Y arm green tuned | Y arm green: Nothing much was disturbed. I touched the steering mirrors and brought GTRY from 0.2 to 0.9.
X arm green: The PDH lock was not very stable mostly because of the low power in green. I changed the oven temperature for the doubler to 36.4 corresponding to maximum green power. GTRX increased from 0.1 to 0.9
Both the X and Y arm green alignment were tuned on the PSL table to their respective beat PDs.
The PSL green shutter was not responding to the medm buttons. I found the PSL green shutter set to 'local' and 'N.O' (these are switches in the shutter controller). I do not see any elog and not sure as to why the controller was even touched in the first place. I set the shutter controls to 'remote' and 'N.C'. |
9620
|
Mon Feb 10 19:56:10 2014 |
rana | Update | ASC | PRM ASC better, but not great yet | Ignoring the OSEM damping loops, the oplev servo loops make it so that the POP ASC loops do not see a simple pendulum plant, but instead see the closed loop response. Since the filter in the OL bank is proportional to f, this means that the open loop gain (OLG):

Which means that the CLG that the ASC sees is going to dip below unity in the band where the OL is on. For example, if the OL loop has a UGF of 5 Hz, it also has a lower UGF of ~0.15 Hz, which means that the ASC needs to know about this modified plant in this band.
For i/eLIGO, we dealt with this in this way: anti-OL in iLIGO |
9619
|
Mon Feb 10 18:59:25 2014 |
Jenne | Update | ASC | PRM ASC better, but not great yet | I have turned off the 3.2Hz res gains in the PRC ASC loops, since those seem to make the loops unstable.
Right now the pitch gain is -0.001, with FM1,3,9 on. Yaw gain is -0.004, with FM1,3,9 on.
Pitch gain can't increase by factor of 2 without oscillating.
I tried to take transfer functions, but I think the ASC situation is really confusing, since I have OSEM damping, oplev damping, and this POP QPD damping on the PRM. It's hard to get coherence without knocking the PRC out of lock, and it keeps looking like my gain is 0dB, with a phase of 0 degrees, from ~1 Hz to ~10 Hz. Outside that range I haven't gotten any coherence. Moral of the story is, I'm kind of puzzled.
Anyhow, as it is right now, the ASC helps a bit, but not a whole lot. I increased the trigger ON value, so that it shouldn't kick the PRM so much. I wish that I had implemented a delay in the trigger, but I'm not in the mood to mess with the simulink diagrams right now. |
9618
|
Mon Feb 10 18:03:41 2014 |
jamie | Update | CDS | 12 core c1sus replacement | I have configured one of the spare Supermicro X8DTU-F chassis as a dual-CPU, 12-core CDS front end machine. This is meant to be a replacement for c1sus. The extra cores are so we can split up c1rfm and reduce the over-cycle problems we've been seeing related to RFM IPC delays.
I pulled the machine fresh out of the box, and installed the second CPU and additional memory that Steve purchased. The machine seems to be working fine. After assigning it a temporary IP address, I can boot it from the front-end boot server on the martian network. It comes up cleanly with both CPUs recognized, and /proc/cpustat showing all 12 cores, and free showing 12 GB memory.
The plan is:
- pull the old c1sus machine from the rack
- pull OneStop, Dolphin, RFM cards from c1sus chassis
- installed OneStop, Dolphin, RFM cards into new c1sus
- install new c1sus back in rack
- power everything on and have it start back up with no problems
Obviously the when of all this needs to be done when it won't interfere with locking work. fwiw, I am around tomorrow (Tuesday, 2/11), but will likely be leaving for LHO on Wednesday. |
9617
|
Mon Feb 10 16:20:28 2014 |
Jenne | Update | General | PRM QPD recentered | In an effort to stop the PRC from wiggling around so much, I recentered the POP QPD after maximizing the POPDC power when locked on carrier. The beam was basically off the QPD in yaw, and at half-range in pitch. |
9616
|
Mon Feb 10 16:02:12 2014 |
Jenne | Update | General | PRM oplev clipped in vacuum | The PRC (locked on carrier so far today), is pretty wobbly. It'll stay locked on carrier, but it's wobbling. The ASC was over-ridden during the vent. While I was looking around for that, I noticed that the PRM oplev sum is very low.
I went into the lab, and turned off / blocked all oplev beams except the PRM beam. I can't tell what it's clipping on, but there is definitely some red glow in the BS chamber (not as much as the stuff that's coming from the ITMY or SRM oplev hitting a tip tilt suspension - that giant spot went away when I turned off the ITMY/SRM oplev laser). The beam going into the vacuum is nice and strong, but the beam coming out is very weak, and has a horizontal line of scatter through it, like it's clipped somewhere in pitch. The PRM oplev sum is currently ~150 cts, when it should be closer to 2,000.

So far, this seems to be livable, but it's definitely disappointing. |
9615
|
Mon Feb 10 09:45:53 2014 |
Steve | Update | VAC | vacuum normal is reached |
Quote: |
Pumpdown completed. IR shutter opened at P1 1 mTorr The block is still in the beam path.
Remember to protect MCR pd before crack up the PSL power.
The ion pump gate valves were just closed by cc1 triggered interlock
The cry pump was "regenerated" during the vent and it's outgassing rate minimized.
CC3 cold cathode gauge was replaced.
Valve configuration for week end:
1, VA6 disconnected to avoid accidental venting the IFO through the annulos
2, VC2 disconnected to insure that the cryo stays closed
3, RGA is not running, It's pressure limit 1e-5 Torr
|
Valve configuration: Vacuum Normal is reached in really 4 days if we do not count overnight rest of roughing.
VA6 and VC2 are reconnected. I'm turning on the RGA next
All 4 ion pumps were vented with air and pumped down to ~ 1e-4 Torr
Ion pumps gate valve control cables are connected and their pumps are disconnected.
The black relay box was tested repeatedly and it stopped misbehaiving.
We were at atmosphere for 13 days. Chamber BS, ITMX, ITMY and ETMY were opened.
Al foil "cups" were placed on the back side OSEMs of PRM. |
Attachment 1: pd77m5d.png
|
|
Attachment 2: 13daysAtm.png
|
|
9614
|
Sat Feb 8 15:14:18 2014 |
ericq | Update | LSC | PRM Sideband Splitting | [ericq]
Today, I kicked the PRM to see the sideband splitting in POP110.
First, we can qualitatively see we moved in the right direction! (See ELOG 9490)

I fit the middle three peaks to a sum of two Lorentzian profiles ( I couldn't get Airy peaks to work... but maybe this is ok since I'm just going to use the location parameter?), and looked at the sideband splitting as a fraction of the FSR, in the same way as in Gabriele's ELOG linked above.
This gave: c / (4 * f55) * (dPhi / FSR) = 0.014 +- .001
Since the PRC length with simultaneous resonance (to 1mm) is given by c / (4 * f11) = 6.773, this means our length is either 6.759m or 6.787m (+- .001). Given the measurement in ELOG 9588, I assume that we are on the short side of the simultaneous resonance. Thus
The sideband splitting observed from this kick indicates a PRC length of 6.759m +- 1mm |
9613
|
Fri Feb 7 17:52:41 2014 |
Koji | Summary | General | Some cleaning up |
- Adjusted the PMC alignment
- Adjusted the IMC length offsets for the MC servo and the FSS servo
- Adjusted the MC alignment. Ran
/opt/rtcds/caltech/c1/scripts/MC/WFS/WFS_FilterBank_offsets
- The Yarm servo was oscillating. Reduced the servo gain from 0.2 to 0.08.
- AS Camera & AS port alignment was adjusted. Now the spot is at the center of the AS camera.
- Cleaned up the ASC offsets on the suspensions (i.e. C1:SUS-***_(PIT|YAW)_OFFSET) by replacing them by the BIAS adjustment.
- Saved the alignment values
- Locked the PRMI with SB with REFL55I for PRCL and AS55Q for MICH
- Aligned the PRM, then checked the alignment of the REFL PDs
- The best POP110I was 500cnt.
- Found REFL165 output was disconnected. It is now restored.
- Used the LSC lockins to figure out the demod phases and the signal amplitudes (relative)
PRCL: 100cnt -> PRM 567.01Hz
Signal in demod Q ch were minimized
REFL11I -19.2deg demod I, Lockin I out (C1:CAL-SENSMAT_PRCL_REFL11_I_I_OUTPUT) 12.6 cnt
REFL33I +130.4deg 1.70cnt
REFL55I +17.0deg 2.30cnt
REFL165I -160.5deg 27.8cnt
MICH: 1000cnt -> ITMX(-1) & (ITMY +1.015 => Minimized the signal in REFL11I to obtain pure MICH)
REFL11I +0.0 REFL11Q +0.119
REFL33I +0.023 REFL33Q -0.012
REFL55I +0.023 REFL55Q -0.113
REFL165I +0.68 REFL165Q +0.038
It seems that REFL165 has almost completely degenerated PRCL and MICH.
- Try to replace ITMX/Y with BS (+0.16) / PRM (-0.084)
ITMX(-1)/ITMY(+1.015) actuation was cancelled by BS (+0.16). This introduces PRCL in REFL11I. This was cancelled by PRM (-0.084)
REFL11I +0.0 REFL11Q +0.13
REFL33I -0.012 REFL33Q 0.025
REFL55I +0.041 REFL55Q -0.45
REFL165I +0.69 REFL165Q +/-0.02
Again, It seems that REFL165 has almost completely degenerated PRCL and MICH.
Locking info:
PRCL:
[Signal source] REFL11I (-19.2deg) x 0.16, OR REFL33I (+130.4deg) x 2.0, OR REFL55I (+17.0deg) x1.0, OR REFL165I (-160.5deg) x0.05
[Trigger] POP110I(-81deg) 100/10, FM trigger 35/2, delay 0.5sec FM2/3/6/9
[Servo] FM4/5 always on. G=-0.02, Limitter ON
[Output] PRM x+1.00
MICH:
[Signal source] AS55Q (-5.5deg) x 1, OR REFL11Q x 0.25, OR REFL55Q x-0.06
[Trigger] POP110I 100/10, FM trigger 35/2, delay 5sec FM2/3/9
[Servo] FM4/5 always on. G=-10, Limitter ON
[Output] ITMX -1.0 / ITMY +1.0 (or +1.015), OR PRM -0.084 / BS +0.16
|
9612
|
Fri Feb 7 13:30:25 2014 |
ericq | Update | General | In-Vac Alignment |
Quote: |
Nice work!! As with all the other RF PDs, POP110's phase likely needs tuning. You want POP110 (and POP22) I-quadratures to be maximally positive when you're locked on sidebands, and maximally negative when locked on carrier. What you can do to get close is lock PRC on carrier, then rotate the POP phases until you get maximally negative numbers. Then, when locked on sideband, you can tweak the phases a little, if need be.
|
Adjusted the angles as Jenne suggested:
- POP22: 105 -> 163
- POP110: 150 -> -83
|
9611
|
Fri Feb 7 13:30:22 2014 |
Gabriele | Update | General | In-Vac Alignment |
Quote: |
Quote: |
I then used the same settings as in ELOG 9554, except I used -1s instead of +1s for the POP110I trigger matrix elements. (I'm not sure why this is different, but I noticed that the PRC would lock on carrier with positive entries here, so I figured we wanted the peaks with opposite sign).
|
Nice work!! As with all the other RF PDs, POP110's phase likely needs tuning. You want POP110 (and POP22) I-quadratures to be maximally positive when you're locked on sidebands, and maximally negative when locked on carrier. What you can do to get close is lock PRC on carrier, then rotate the POP phases until you get maximally negative numbers. Then, when locked on sideband, you can tweak the phases a little, if need be.
|
Very good news! We should have a look at the POP110 sideband peak splitting, to see if we really got the right PRC length... |
9610
|
Fri Feb 7 13:00:52 2014 |
Jenne | Update | General | In-Vac Alignment |
Quote: |
I then used the same settings as in ELOG 9554, except I used -1s instead of +1s for the POP110I trigger matrix elements. (I'm not sure why this is different, but I noticed that the PRC would lock on carrier with positive entries here, so I figured we wanted the peaks with opposite sign).
|
Nice work!! As with all the other RF PDs, POP110's phase likely needs tuning. You want POP110 (and POP22) I-quadratures to be maximally positive when you're locked on sidebands, and maximally negative when locked on carrier. What you can do to get close is lock PRC on carrier, then rotate the POP phases until you get maximally negative numbers. Then, when locked on sideband, you can tweak the phases a little, if need be. |
9609
|
Fri Feb 7 12:43:12 2014 |
ericq | Update | General | In-Vac Alignment | [ericq]
PRC Locked on Sidebands 
Jenne reminded me that if we change a cavity, phases can change... So, first, I locked the PRC on the carrier, and then gave it MICH and PRCL excitations to optimize the AS55 and REFL55 phase rotation angles by looking at the excitation demodulated outputs of the unused quadrature (i.e. we want all of MICH to be in AS55 Q, so I rotated the phase until C1:CAL-SENSMAT_MICH_AS55_I_I_OUTPUT was zero on average).
This resulted in:
- AS55: 7 -> -5.5
- REFL55: 73 -> 86.9
I then used the same settings as in ELOG 9554, except I used -1s instead of +1s for the POP110I trigger matrix elements. (I'm not sure why this is different, but I noticed that the PRC would lock on carrier with positive entries here, so I figured we wanted the peaks with opposite sign).
So far, it seems more stable than when we were doing the demodulation phase measurements, it's been locked for >15 minutes without me having to tweak the gains or the alignment from the carrier locked case. |
9608
|
Thu Feb 6 16:41:31 2014 |
ericq | Update | General | In-Vac Alignment | [Manasa, ericq]
Both arms have been aligned via ASS. PRC locked on carrier.
SB locking hasn't happened yet...
Details:
- Aligned MC at low power, measured spots
- Found ITMX AS, REFL spots on cameras. Couldn't find ITMY spot. Found x-arm flashes with PRM aligned.
- MC Refl Y1 mirror was replaced with the 90:10 BS; blocked WFS path until MC was aligned
- PMC power was increased (~1.4W directly before PSL shutter)
- Touched up MC alignment, reactivated high power autolocker, measured spots.
- Locked x-arm on IR, ran ASS
- Found ITMY spot on AS camera, locked y-arm, ran ASS on both arms
- Checked green beams. Y-arm locks at around .6, X-arm at around .2 (input steering needs adjustment)
- Centered beams on WFS, reset filter bank offsets, turned on WFS
- Aligned PRM, locked PRC on carrier.
- Tried locking PRC on sideband, had troubles. Looks like there is some increased seismic activity; might be the culprit.
|
9607
|
Thu Feb 6 11:14:07 2014 |
Steve | Update | VAC | pumpdown completed | Pumpdown completed. IR shutter opened at P1 1 mTorr The block is still in the beam path.
Remember to protect MCR pd before crack up the PSL power.
The ion pump gate valves were just closed by cc1 triggered interlock
The cry pump was "regenerated" during the vent and it's outgassing rate minimized.
CC3 cold cathode gauge was replaced.
Valve configuration for week end:
1, VA6 disconnected to avoid accidental venting the IFO through the annulos
2, VC2 disconnected to insure that the cryo stays closed
3, RGA is not running, It's pressure limit 1e-5 Torr
|
Attachment 1: pd77.png
|
|
9606
|
Wed Feb 5 20:41:57 2014 |
Den | Update | LSC | calibrated spetra from OAF test | We did online adaptive filtering test with IMC and arms 1 year ago (log 7771). In the 40m presentations I can still see the plot with uncalibrated control spectra that was attached to that log. Now it the time to attach the calibrated one.
Template is in the /users/den/oaf |
Attachment 1: oaf_cal.pdf
|
|
9605
|
Wed Feb 5 19:53:51 2014 |
Steve | Update | VAC | pumpdown stops for the day at 14 Torr |
|
Attachment 1: stopHere.png
|
|
9604
|
Wed Feb 5 19:36:50 2014 |
Steve | Update | VAC | pumpdown at 25 Torr |
Quote: |
[Steve, Manasa]
I checked the alignment one last time. The arms locked, PRM aligned, oplevs centered.
We went ahead and put the heavy doors ON. Steve is pumping down now!
|
The ion pumps were vented just before pumpdown and their gate valves were opened.
This is an effort to minimize a possible leak through their gates.
Is there a volunteer who goes home late and would close off the roughing? tonight |
Attachment 1: pd77at3.5h.png
|
|
9603
|
Wed Feb 5 18:36:56 2014 |
rana | Frogs | elog | MicroSoft BingBot is attacking us | The ELOG was frozen, with this in the .log file:
GET /40m/?id=1279&select=1&rsort=Type HTTP/1.1
Cache-Control: no-cache
Connection: Keep-Alive
Pragma: no-cache
Accept: */*
Accept-Encoding: gzip, deflate
From: bingbot(at)microsoft.com
Host: nodus.ligo.caltech.edu
User-Agent: Mozilla/5.0 (compatible; bingbot/2.0; +http://www.bing.com/bingbot.htm)
(hopefully there's a way to hide from the Bing Bot like we did from the Google bot)
|
9602
|
Wed Feb 5 15:39:41 2014 |
manasa | Update | General | We are pumping down | [Steve, Manasa]
I checked the alignment one last time. The arms locked, PRM aligned, oplevs centered.
We went ahead and put the heavy doors ON. Steve is pumping down now! |
Attachment 1: pre_pump_down.png
|
|
9601
|
Wed Feb 5 15:26:57 2014 |
manasa | Update | General | Ready for pump down?? |
Quote: |
This sounds great! The only suggestion that I have is for checking POP. If you have the beam on the camera, you can hold a card in front of each mirror, and find out where the edge of the beam is. Introduce the card from the side, and watch for the point where you just start to see the beam on the camera be obstructed. Repeat for the other side, and you have an idea of the centering of the beam.
I think this is most important for the in-vac mirrors, since the beam is large-ish, and we have to hit both steering mirrors at ~45 degrees.
|
[EricQ, Manasa]
This check was done and we had to move one of the steering mirrors in pitch. Else, everything was just fine.
In-vacuum pictures of PR2 and PR3 new positions were taken. MC spot positions measured to be < 1mm and oplevs were centered. |
9600
|
Wed Feb 5 09:28:32 2014 |
Steve | Update | SUS | ETMY damping restored | ETMY damping restored. |
9599
|
Wed Feb 5 08:14:59 2014 |
Steve | Update | VAC | vacuum computers are back without safety |
Quote: |
1, Reset works
2, Default values are lost. They actually reversed into open and turn on everything when power recycle c1vac1, c1vac2 or 24V dc power supply. This can vent the IFO in an event of power failure!
It's may be the time to go back to an isolated, vacuum controller only computer.
The present valve configuration at Atm1
a, all annuloses are vented
b, valve cables disconnected at: VC1 and 4 IP gate valves
c, the RGA is off and it is pumped by the Maglev through VM2
d, cryo pump is being roughed with TP3 It's outgassing rate was 25 mTorr / min
e, Roughing hose is disconnected
|
Vacuum checks before pump down:
1, check intermittent contact issue in black relay box
2, reconnect cable to ion pumps and vent them to atm before pump down
3, repeat power shutdown-reboot at atmosphere |
9598
|
Tue Feb 4 23:01:24 2014 |
Jenne | Update | General | Ready for pump down?? |
This sounds great! The only suggestion that I have is for checking POP. If you have the beam on the camera, you can hold a card in front of each mirror, and find out where the edge of the beam is. Introduce the card from the side, and watch for the point where you just start to see the beam on the camera be obstructed. Repeat for the other side, and you have an idea of the centering of the beam.
I think this is most important for the in-vac mirrors, since the beam is large-ish, and we have to hit both steering mirrors at ~45 degrees. |
9597
|
Tue Feb 4 17:40:19 2014 |
manasa | Update | General | Ready for pump down?? |
Quote: |
* PRM suspension has not been behaving well. PRM is being kicked around every 5-10 seconds when the PRC is aligned (as seen on REFL camera). We are not sure where this is coming from. The first time we saw this happening was when we were trying to lock PRC at low power even before we took the heavy doors off. So we are pretty sure this is not caused by the foil cover on the OSEMs. We tried turning ON/OFF the oplev servo, turning ON/OFF the damping loops and also checked the connections in the feedthrough and satellite box for the PRM. The OSEM sensor values for the suspension also seem to match the ones on the wiki.
|
This is solved.
The ASC for PRC for left turned ON. Turning it OFF solved the problem.
If there is no feedback regarding the POP alignment or anything to check with modified PRC length, we will close tomorrow morning. |
9596
|
Tue Feb 4 16:43:50 2014 |
manasa | Update | General | IFO alignment update | [EricQ, Manasa]
We are close to the end of the vent except for a couple of issues.
* POP is not visible on the IR card. But we see POP flashes unclipped on the camera and also spikes in POP DC. So we are assuming that the POP path hasn't gone far off. If anybody has suggestions for a better method to check this, we could give it a try.
* PRM suspension has not been behaving well. PRM is being kicked around every 5-10 seconds when the PRC is aligned (as seen on REFL camera). We are not sure where this is coming from. The first time we saw this happening was when we were trying to lock PRC at low power even before we took the heavy doors off. So we are pretty sure this is not caused by the foil cover on the OSEMs. We tried turning ON/OFF the oplev servo, turning ON/OFF the damping loops and also checked the connections in the feedthrough and satellite box for the PRM. The OSEM sensor values for the suspension also seem to match the ones on the wiki.
Closeup checklist
Center beam on all AS optics
-
GET CAMERA IMAGES OF EVERYTHING
- We must get images right before closing, right after closing, etc.
Make sure REFL is clear
dither PRM, see motion on AP tables
Make sure AS is clear
dither BS/ITM, see motion on AP tables
Using IPANG/POS pick-off mirrors, center beams on:
IPPOS
IPANG (IPANG aligned to be low in pitch. The beam comes out of vacuum unclipped and hits the first steering mirror outside vacuum at its lower edge)
- Check green alignment in the arms and make sure the transmitted green reaches the PSL table.
-
Check all OpLevs centered, in and out of vacuum
-
Close PSL shutter & green shutters at the ends
- Check jam nuts
|
9595
|
Tue Feb 4 01:02:03 2014 |
Koji | Update | Green Locking | ETMX green power | Manasa, Steve: Please revisit the Xend oven temperature again.
I found that the X end SLOW control was left on for ~15days. The output of the filter had grown to ~2e7.
This yielded the laser temperature pulled with the maximum output of the DAC.
This was the cause of the power reduction of the X end SHG; phase matching condition was changes as the wavelength of the IR was changed.
Once the SLOW output was reset, the green REFL was reduced from 4000cnt to 1800cnt. |
Attachment 1: Screenshot-Untitled_Window.png
|
|
9594
|
Tue Feb 4 00:42:18 2014 |
Koji | Update | General | X arm aligned for IR/GR | The X arm was also aligned for the IR by hand and ASS. Also the X end green PZT was aligned to make the TEM00 mode reasonably locked.
What I did:
- Looked at the ITMXF camera. It seemed that the green beam was hitting the mirror.
- Went to the end. Looked at the X end green REFL. Tuned coarse alignment of the ETMX so that the beam was (retro-)reflected to the Faraday and the REFL PD.
- Looked at the ETMX face from the view port. Tried to locate the spot from the ITMX by shaking the ITMX alignment with 0.1 and then 0.01 increments.
- After some struggle with the ETMX and ITMX alignment, resonant fringes were found on the ETMY face while I still looked at the ETMX.
- Once the ITMX/ETMX were aligned, the BS needed to be aligned. But of course there was no IR fringe.
- Returned to the original alignment of the ITMX to find the ITMX spot on the AS camera.
Then gradually moved the ITMX to the aligned value for the green while tracking the michelson alignment with the BS.
This made the AS spots at the upper left edge of the AS video image.
- This was enough to find the IR spikes at TRX. Then the ETMX was touched to maximize the transmission.
- Lock the cavity. Use the ASS to optimize the alignement.
- Once the arm mirrors were aligned, the Xend PZT was also adjusted to have TEM00 for the green beam.
Now I leave the IFO with ITMX/Y, ETMX/Y and BS aligned. As I wrote above, the AS spot is very high at the AS camera.
We need to revisit the AS steering (SR TTs?) to ensure the AS beam unclipped. |
9593
|
Mon Feb 3 23:31:33 2014 |
Manasa | Update | General | Alignment update / Y arm locked | [EricQ, Manasa, Koji]
We measured the spot positions on the MC mirrors and redid the MC alignment by only touching the MC mirror sliders. Now all the MC spots are <1mm away from the center.
We opened the ITMY and ETMY chambers to align the green to the arm. The green was already centered on the ITMY. We went back and forth to recenter the green on the ETMY and ITMY (This was done by moving the test masses in pitch and yaw only without touching the green pointing) until we saw green flashes in higher order modes. At this point we found the IR was also centered on the ETMY and a little low in pitch on ITMY. But we could see IR flashes on the ITMYF camera. We put back the light doors and did the rest of the alignment using the pitch and yaw sliders.
When the flashes were as high as 0.05, we started seeing small lock stretches. Playing around with the gain and tweaking the alignment, we could lock the Y arm in TEM00 for IR and also run the ASS. The green also locked to the arm in 00 mode at this point. We aligned the BS to get a good AS view on the camera. ITMX was tweaked to get good michelson. |
9592
|
Mon Feb 3 15:57:52 2014 |
Steve | Update | Green Locking | ETMX green power | There was 0.2 mW green at the X end.
The doubling oven temp was changed from 37.5 to 36 degrees C
Power at green shutter 3 mW The alignment was not touched. |
Attachment 1: ETMXgreen.jpg
|
|
Attachment 2: XgreenOven36C.jpg
|
|
9591
|
Mon Feb 3 10:17:14 2014 |
Steve | Update | PEM | dusty surfaces | Please wet WIPE before opening chamber or optical table ! !
with methanol soaked kimwipes.
The Met One particle counter is located on CES wall, just behind ITMX chamber.
The numbers are not so bad, but have you ( ...a) asked the IFO lately? |
Attachment 1: AP-ISCT.jpg
|
|
Attachment 2: BSC-rim.jpg
|
|
Attachment 3: ITMX-ISCT.jpg
|
|
Attachment 4: 12daysAtm.png
|
|
9590
|
Fri Jan 31 19:29:36 2014 |
Gabriele | Summary | LSC | PRC and SRC lengths | Today we measured the missing distance to reconstruct SRC length.
I also changed the way the mirror positions are reconstructed. In total for PRC and SRC we took 13 measurements between different points. The script runs a global fit to these distances based on eight distances and four incidence angles on PR2, PR2, SR2 and SR3. The optimal values are those that minimize the maximum error of the 13 measurements with respect to the ones reconstructed on the base of the parameters. The new script is attached (sorry, the code is not the cleanest one I ever wrote...)
The reconstructed distances are:
Reconstructed lengths [mm]:
LX = 6771
LY = 6734
LPRC = 6752
LX-LY = 37
LSX = 5493
LSY = 5456
LSRC = 5474
The angles of incidence of the beam on the mirrors are very close to those coming from the CAD drawing (within 0.15 degrees):
Reconstructed angles [deg]:
aoi PR3 = 41.11 (CAD 41)
aoi PR2 = 1.48 (CAD 1.5)
aoi SR3 = 43.90 (CAD 44)
aoi SR2 = 5.64 (CAD 5.5)
The errors in the measured distances w.r.t. the reconstructed one are all smaller than 1.5 mm. This seems a good check of the global consistency of the measurement and of the reconstruction method.
NOTES: in the reconstruction, the BS is assumed to be exactly at 45 degrees; wedges are not considered.
|
Attachment 1: map_jan31st.pdf
|
|
Attachment 2: survey_v3.zip
|
9589
|
Fri Jan 31 18:41:25 2014 |
manasa | Update | General | IFO alignment update | [EricQ, Gabriele, Manasa]
We found we had lost the Y arm pointing from yesterday. We tried to recover the pointing for a couple of hours and finally decided to take the ETMY heavy door off.
The input beam was aligned to the Y arm. We also got AS and REFL out of vacuum and on the cameras.
We put back the light doors and tried to lock the arms, but did not succeed as yet.
Things to do:
1. Lock arms for IR
2. Realign POP path
3. Recenter all oplevs
4. Try to check the state of PRC after the length change
5. Take in-vacuum pictures |
9588
|
Thu Jan 30 19:00:25 2014 |
Gabriele | Summary | LSC | PRC length changed | [Manasa, EricQ, Gabriele]
Today we changed the PRC length translating PR2 by 27 mm in the direction of the corner. After this movement we had to realign the PRC cavity to get the beam centered on PRM, PR2, PR3, BS (with apertures) and ITMY (with aperture). To realign we had to move a bit both PR2 and PR3. We could also see some flashes back from the ETMY . //Edit by Manasa : We could see the ETMY reflection close to the center of the ITMY but the arm is not aligned or flashing as yet//.
After the realignment we measured again the PRC length with the same method of yesterday. We only had to change one of the length to measure, because it was no more accessible today. The new map is attached as well as the new script (the script contains also the SRC length estimation, with random numbers in it).
The new PRC length is 6753 mm, which is exactly our target!
The consistency checks are within 5 mm, which is not bad.
We also measured some distances to estimate the SRC length, but right now I'm a bit confused looking at the notes and it seems there is one missing distance (number 1 in the notes). We'll have to check it again tomorrow.
|
Attachment 1: map_jan30.pdf
|
|
Attachment 2: survey_v3_jan30.m
|
clear all
global sos_lx sos_ly sos_cx sos_cy tt_lx ...
tt_ly tt_cx tt_cy sos_sx sos_sy sos_dy
%% Survey of the PRC+SRC lengths %%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%% measured distances
d_MB2_MY = 2114 + 27 + 9;
... 446 more lines ...
|
9587
|
Thu Jan 30 11:59:03 2014 |
manasa | Update | CDS | fb timing was off |
Quote: |
Since this morning, the fb's timing has been off. Steve pointed it out to me earlier today, but I didn't have a chance to look at it until now.
This was different from the more common problem of the mx stream needing to be restarted - that causes 3 red blocks per core, on all cores on a computer, but it doesn't have to be every computer. This was only one red block per core in the CDS FE status screen, but it was on every core on every computer.
The error message, when you click into the details of a single core, was 0x4000. I elog searched for that, and found elog 6920, which says that this is a timing issue with the frame builder. Since Jamie had already set things on nodus' config correctly, all I did was reconnect the fb to the ntp:
fb$ sudo /etc/init.d/ntp-client restart
As in elog 6920, the daqd stopped, then restarted itself, and cleared the error message. It looks like everything is good again.
I suspect (without proof) that this may have to do with the campus network being down this morning, so the computers couldn't sync up with the outside world.
|
The above timing problem has been repeating (a couple of times this week so far). It does not seem to be related to the campus network.
The same solution was applied. |
9586
|
Wed Jan 29 21:01:03 2014 |
Gabriele | Summary | LSC | PRC length measurement analysis | [Manasa, EricQ, Gabriele]
We managed to measure the PRC length using a procedure close to the one described in slog 9573.
We had to modify a bit the reference points, since some of them were not accessible. The distances between points into the BS chamber were measured using a ruler. The distances between points on different chambers were measured using the Leica measurement tool. In total we measured five distances, shown in green in the attached map.
We also measured three additional distances that we used to cross check the results. These are shown in the map in magenta.
The values of the optical lengths we measured are:
LX = 6828.96 mm
LY = 6791.74 mm
LPRC = 6810.35 mm
LX-LY = 37.2196 mm
The three reference distances are computed by the script and they match well the measured one, within half centimeter:
M32_MP1 = 117.929 mm (measured = 119 mm)
MP2_MB3 = 242.221 mm (measured = 249 mm)
M23_MX1p = 220.442 mm (measured = 226 mm)
See the attached map to see what the names correspond to.
The nominal PRC length (the one that makes SB resonant without arms) can be computed from the IMC length and it is 6777 mm. So, the power recycling cavity is 33 mm too long w.r.t. the nominal length. This is in good agreement with the estimate we got with the SB splitting method (4cm).
According to the simulation in the wiki page the length we want to have the SB resonate when the arms are there is 6753 mm. So the cavity is 57 mm too long.
Attached the new version of the script used for the computation. |
Attachment 1: map.pdf
|
|
Attachment 2: script.zip
|
9585
|
Wed Jan 29 16:36:37 2014 |
Koji | Summary | General | High power beam blasting of the aLIGO RFPD | [Rich, Jay, Koji]
We blasted the aLIGO RF PD with a 1W IR beam. We did not find any obvious damage.
Rich and Jay brought the PD back to Downs to find any deterioration of the performance with careful tests.
The power modulation setup is at the rejection side of the PBS in front of the laser source.
I checked the beams are nicely damped.
As they may come back here tomorrow, a power supply and a scope is still at the MC side of the PSL enclosure. |
9584
|
Tue Jan 28 23:32:12 2014 |
Koji | Update | General | X/Y arm locked with the IR beam | [Koji EricQ]
The both arms have been locked with IR and aligned by ASS.
The IFO was left with ITMX/Y, ETMX/Y, BS, and PRM aligned, and the PSL shutter closed.
YARM
SIGNAL PATH: POY11I(+45dB)->YARM(G=+1.0)->ETMY
NORM: TRYx10
TRIG: TRY 0.01up/0.005down
FM TRIG: FM2/3/6/7/8/9 0.01up/0.05down, 0.5 sec delay
XARM
SIGNAL PATH: POX11I(+45dB)->XARM(G=+4.0)->ETMX
NORM: TRXx10
TRIG: TRX 0.01up/0.005down
FM TRIG: FM2/3/6/7/8/9 0.01up/0.05down, 0.5 sec delay
For decent locks, it was necessary that the offset of the error signals are trimmed at the input filters
even after running LSCoffset.py script.
Once the cavities were aligned for the IR, we could see the green beams are also flashing.
The Y arm was actually locked with the green with a TEM00 mode |
Attachment 1: good_alignment.png
|
|
9583
|
Tue Jan 28 22:24:46 2014 |
ericq | Update | General | Further Alignment | [Masasa, ericq]
Having no luck doing things remotely, we went into the ITMX chamber and roughly aligned the IR beam. Using the little sliding alignment target, we moved the BS to get the IR beam centered on ITMX, then moved ITMX to get good michelson fringes with ITMY. Using an IR card, found the retroflection and moved ETMX to make it overlap with the beam transmitted through the ITM. With the PRM flashing, X-arm cavity flashes could be seen. So, at that point, both the y-arm and x-arm were flashing low order modes. |
9582
|
Tue Jan 28 16:26:40 2014 |
Steve | Update | VAC | vacuum computers are back without safety |
Quote: |
[Steve Koji]
We pushed the reset button of c1vac1 and c1vac2 and the vacuum screen is back.
First, we pushed the reset button of c1vac1 and pushed the one on c1vac2.
This did not bring c1vac2 up. We pushed the reset of c1vac2 again and now everything of the vacuum screen is back.
|
1, Reset works
2, Default values are lost. They actually reversed into open and turn on everything when power recycle c1vac1, c1vac2 or 24V dc power supply. This can vent the IFO in an event of power failure!
It's may be the time to go back to an isolated, vacuum controller only computer.
The present valve configuration at Atm1
a, all annuloses are vented
b, valve cables disconnected at: VC1 and 4 IP gate valves
c, the RGA is off and it is pumped by the Maglev through VM2
d, cryo pump is being roughed with TP3 It's outgassing rate was 25 mTorr / min
e, Roughing hose is disconnected |
Attachment 1: c1vac1_vac2areback.png
|
|
9581
|
Tue Jan 28 11:13:50 2014 |
Koji | Update | VAC | vacuum monitor is still blank | [Steve Koji]
We pushed the reset button of c1vac1 and c1vac2 and the vacuum screen is back.
First, we pushed the reset button of c1vac1 and pushed the one on c1vac2.
This did not bring c1vac2 up. We pushed the reset of c1vac2 again and now everything of the vacuum screen is back. |
9580
|
Tue Jan 28 09:51:47 2014 |
Steve | Update | IOO | low power pointing |
PSL output is stable. |
Attachment 1: IOO4dlowpr.png
|
|
9579
|
Mon Jan 27 21:36:35 2014 |
ericq | Update | General | Vent so far | After turning the slow FSS threshold down, the mode cleaner stays locked enough to do other things. We were able to align the tip tilts to the y-arm such that we were able to get some flashes in what looks like a TM00-ish mode. (It was necessary to align the PRM such that there was some extra power circulating in the PRC to be able to see the IR flashes on the ITMY face camera) This is enough to convince us that we are at least near a reasonable alignment, even though we couldn't lock to the mode.
The x-arm was in a hairier situation; since the green beam wouldn't flash into any modes, we don't even know that a good cavity axis exists. So, I used the green input PZTs to shine the green beam directly on the earthquake stops on the ITMX cage, and then inferred the PZT coordinates that would place the green beam roughly on the center of ITMX. I moved the ETMX face camera such that it points at the ETMX baffle. I tried looking for the retroreflected green spot to no avail. Hopefully tomorrow, we can get ourselves to a reasonably aligned state, so we can begin measuring the macroscopic PRC length. |
9578
|
Mon Jan 27 17:49:46 2014 |
rana | HowTo | Computer Scripts / Programs | sendmail started on nodus: fixing SwiftMail on Dokuwiki | Since the recent filesystem fracas, the new accounts could not be created on nodus / dokuwiki (for the controls workshop, for example).
I started sendmail on nodus using the command: sudo /etc/init.d/sendmail start
and the SwiftMail plugin on there is now sending out the confirmation emails again. This will happen each time we reboot nodus, so let's replace it. |
|